OpenStudy (anonymous):

Solve for x and y: 5/4x-3/4y=9/4 5/3y-7/3x=15/3

OpenStudy (anonymous):

i got to 5x-3y=9 and -7x+5y=15

OpenStudy (anonymous):

Okay, so we can use elimination: This one is gonna be interesting though. If we multiply equation 1 by 4 and equation 2 by 3, then we get rid of the denominators. 5x-3y=9 -7x+5y=15 as you have... next we multiply equation 1 by 5 and equation 2 by 3. 25x-15y=45 -21x+15y=45 4x=90 x=90/4 Now plug in x into one of the equations and solve for y. Hope this helps. http://www.tutorsean.net Follow me on Facebook: Tutor Sean.